OpenStudy (anonymous):

Doug is a salesperson in a retail store and earns $75 per week, plus 15% of his weekly sales. If Doug makes $555 one week, what are his sales for that week? $3,200 $627 $566 $1,914

OpenStudy (anonymous):

@green_1

OpenStudy (karolynxx):

75 + 15x/100 = 555 then 15x/100 = 555-75 finally x = 480 * 100 / 15 Which is 3200.

OpenStudy (anonymous):

Thank you :)